OpenStudy (anonymous):

Help with simplifying... (10/x) -5 over 5

OpenStudy (anonymous):

So this is the question: \[\frac{ 10 }{ x } - 5 \over 5\]

OpenStudy (anonymous):

Yes

jhonyy9 (jhonyy9):

10 10 5 10 -5x 5(2-x) --- - 5 = ---- - ---- = --------- = --------- x x 1 x x so this is divided again by 5 than what will be 5(2-x) ------ x 5(2-x) 1 2 - x -------- = ------- * ----- = ------- 5 x 5 x